23

24

25

Suppose that you have $14,000 to invest and you are trying to decide between investing in project A or project B. If you invest in project A, you will receive a payment of $16,500 at the end of 2 years. If you invest in project B, you will receive a payment of $25,000 at the end of 11 years. Assume the annual interest rate is 5 percent and that both projects carry no risk. Instructions: Round your answers to the nearest dollar. Do not round intermediate calculations a. What is the present value of each project? Present Value Project A Project B b. Which project would you choose for your investment?
Suppose that your business has an opportunity to buy a machine that produces widgets. The machine can produce 9,000 widgets per year that you can sell for a profit of $2.40 per widget. With proper maintenance, the machine can run indefinitely. You decide that your minimum annual rate of return for any investment is 9 percent. What is the maximum price you would be willing to pay for the machine? Instructions: Round your answer to the nearest dollar. Maximum price-
You are trying to decide if you should rent an ice cream truck to drive around the city park to help pay for your college tuition. Assume that it costs $300 per day to rent the truck and pay the driver's wages. Your marketing data show that if it does not rain, you can expect profits of $180 per day (after paying all expenses including the truck rental and the driver's wages). If it rains, then expected profits fall to $64 per day. Based on historical weather data, you can expect the weather to be rainy 25 percent of the time. What is your expected daily rate of return on the rental cost of the truck and driver? Instructions: Round your answer to 1 decimal place. Rate of return
Answer #1

23.

Present value of project A = -14000 + 16500*(P/F,5%,2) = -14000 + 16500 * 0.907029 = 966

Present value of project B = -14000 + 25000*(P/F,5%,11) = -14000 + 25000 * 0.584679 = 617

Project A should be selected as it has higher NPV

24.

Total profit per year = 9000 * 2.4 = 21600

i = 9% = 0.09

Present value of perpetual annuity = A/ i

Present value of machine (maximum price) = 21600 / 0.09 = 240000

25.

return per day without rain = 180/300 = 60%

return per day with rain = 64/300 = 21.33%

25% days its raining

Avg return = 0.75*0.6 + 0.25*0.21333 = 0.5033 = 50.3%